# Periodic classical trajectories and quantum scars in many-spin systems

### v2

@@ IV.1 Equations of motion
Where necessary, the set of equations (5, 6) is to be solved numerically using the 4th-order Runge-Kutta algorithm with the discretization time steps $\delta t$ being small enough to make the results insensitive to making $\delta t$ smaller; typically $\delta t=0.001$.
@@
where the projections are the solutions of Eqs.(5, 6).
@@ IV.2 Periodic trajectories
Due to the translational invariance of both the Hamiltonian (III) and the initial conditions (2), the classical spins initially polarised in the same direction exhibit the same time evolutions and hence remain parallel to each other at all subsequent moments of time with the direction of their common polarisation evolving as a function of time. As a result, the calculation of a many-spin trajectory $\mathcal{S}(t)$ is reduced to computing the same one-spin trajectory $\mathbf{S}_{i}(t)=\mathbf{S}_{\text{p}}(t)$ for each spin on the lattice. Since the one-spin trajectory $\mathbf{S}_{\text{p}}(t)$ is limited to a two-dimensional manifold - the surface of a unit sphere $|\mathbf{S_{\text{p}}}|=1$, it normally closes onto itself and thus necessarily becomes periodic. The corresponding many-spin phase space trajectory, to be denoted as $\mathcal{S}_{\text{p}}(t)$, also becomes periodic.
@@ IV.3 Librations and rotations
Dependent on the parameters of the Hamiltonian $\mathcal{H}_{\text{p}}$ and on the initial orientation $\mathbf{S}_{\text{p}}(0)$, there are two possible kinds of periodic trajectories $\mathbf{S}_{\text{p}}(t)$, which, by analogy with the description of a pendulum[43], we call “librations” and “rotations”. Librations are realised by a single trajectory on the spherical surface $|\mathbf{S_{\text{p}}}|=1$ connecting all points having the same energy. The rotations are realised by two disconnected trajectories – both corresponding to the same energy. Librations and rotations transition to each other as a function of Hamiltonian parameters and/or initial energy through a separatrix as illustrated in Fig.2.
@@ IV.4 Lyapunov instabilities: general considerations
Practical numerical computation of $\lambda_{\text{max}}$ requires one to perform a large number of resets (contractions) of $|\delta\mathcal{S}(t)|$ as described in Refs.[121, 122] and in Appendix A.
Largest Lyapunov exponents $\lambda_{\text{max}}$ and the entire Lyapunov spectra $\{\lambda_{k}\}$ of classical spin systems have previously been investigated [44, 120] but only in the vicinity of ergodic trajectories, which nonperiodically cover the energy shell of the system. If a trajectory is ergodic, it is supposed to pass in the vicinity of every point on the energy shell, thereby “collecting” all possible local growth rates of $|\delta\mathcal{S}(t)|$ and averaging over them. As a result, $\lambda_{\text{max}}$ and the entire Lyapunov spectrum do not depend on the choice of the initial conditions $|\mathcal{S}(0)|$. The convergence of the procedure for computing $\lambda_{\text{max}}$ is, in fact, a good measure of the ergodisation time of the underlying dynamics[45].
@@
Generically, different wave vectors $\mathbf{q}_{k}$ are supposed to correspond to different values of $\lambda_{k}$ with the exception of symmetry-related or accidental degeneracies. A symmetry-related degeneracy occurs, in particular, when the lattice has inversion symmetry (as is the case for our investigation), and, as a result, pairs of wave vectors ±$\pm\mathbf{q}_{k}$ correspond to the same $\lambda_{k}$. In such a case, a generic Lyapunov exponent is at least double-degenerate, the possible non-degenerate cases correspond to $\mathbf{q}_{k}=0$ and $\mathbf{q}_{k}=\pi/L$.
@@ IV.5.1 Overview
Figure 4: Dependence of periodic Lyapunov exponent $\lambda_{\text{p}}$ on chain length $L$ for three different values of $J$: (a)  $J=0.79$, (b) $J=1.76$ and (c) $J=2.23$. Panel (a) corresponds to librations, while panels (b) and (c) correspond to rotations. Points represent numerically computed $\lambda_{\text{p}}$. Red lines are the fits based on Eq.(12) with parameters indicated in the plot legends. Points marked as “Stable” correspond to the $\lambda_{\text{p}}=0$ verified as explained in Appendix A. In a few cases marked as entering Arnold diffusion regime, Lyapunov instability around a periodic trajectory is followed by anomalously long-living nearly quasiperiodic behavior — see Section V for further discussion.
@@
Let us, finally, mention that the largest Lyapunov exponent $\lambda_{\text{p}}$ for a periodic trajectory can be either larger or smaller than the largest Lyapunov exponent $\lambda_{\text{e}}$ for a nonperiodic ergodic one on the same energy shell. The plots of $\lambda_{\text{e}}$ for different chain lengths and interaction constants can be found in Fig. 23 of Appendix A. For example, $\lambda_{\text{p}}\approx 0.35$ for a chain with $J=1.76$ in the limit of large $L$ [see Fig. 4(b)], which is significantly smaller than the ergodic exponent $\lambda_{\text{e}}\approx 1.13$ for the same $J$ (see Fig. 23). On the other hand, the periodic exponent near the separatrix $J=J^{*}$ reaches the value $\lambda_{\text{p}}\simeq 1.43$ (see Fig. 7), which is noticeably larger than the ergodic exponent $\lambda_{\text{e}}\simeq 0.67$ in Fig. 23 for the same $J$.
@@
Figure 5: Spatial Fourier spectra $f_{\lambda_{\text{p}}}(q)$ of the Lyapunov vectors defined by Eq. (11). The chain lengths $L$ and the coupling constants $J$ are indicated in plot legends. The allowed values of wavenumbers are $q=\frac{2\pi}{L}k$, where $k$ is an integer in the interval $[-L/2,L/2]$.
@@
$f_{\lambda_{\text{p}}}(q)\equiv\frac{1}{L}\left|\sum_{m=1}^{m=L}\delta\mathbf{S}_{m}\ e^{-iqm}\right|^{2}$ (11)
of the Lyapunov vectors $\delta{\cal S}$ corresponding to $\lambda_{\text{p}}$. Here $\delta\mathbf{S}_{m}\equiv(\delta S_{m}^{x},\delta S_{m}^{y},\delta S_{m}^{z})$ are the components of the high-dimensional vector $\delta{\cal S}$ entering Eq.(10); the Fourier amplitude inside the absolute-value bars |||...|| … | is a three-dimensional vector with $x$, $y$ and $z$ complex-valued projections, $|...|^{2}$ is the sum of the squares of the moduli of those projections. The Fourier spectra in Fig. 5 are computed for two chain lengths $L=18$ and $L=20$ and for two interaction constants $J$ representing the libration and the rotation regimes for each length. In all four cases, one can see two dominant Fourier peaks corresponding to ±$\pm q_{\text{p}}$.
The underlying translational symmetry breaking helps us to explain the distinct behavior of $\lambda_{\text{p}}(J)$ plotted in Fig.3: In both panels of Fig.3, there are intervals of $J$ where Lyapunov exponents $\lambda_{\text{p}}$ for two chains of different lengths coincide with each other, and also there are intervals where the two Lyapunov exponents suddenly become different, because one of of the two functions $\lambda_{\text{p}}(J)$ exhibits a kink. The pair of plots in Fig. 3(a) and another pair in Fig. 3(b) – both correspond to chain lengths such that one chain is twice the length of the other. Therefore, the wave numbers $q=\frac{2\pi}{L}n$, where $n$ is an integer and $L$ the length of the shorter chain, are the allowed wave numbers for the both chains, but then the longer chain has additional allowed wave numbers that are not available for the shorter one. Thus, if the largest Lyapunov exponent for the longer chain corresponds to $q$ available for the shorter chain, then the two Lyapunov exponents $\lambda_{\text{p}}$ are equal to each other. If, however, the above $q$ for the longer chain is not available for the shorter one, then the values of $\lambda_{\text{p}}$ become different, with the larger $\lambda_{\text{p}}$ necessarily corresponding to the longer chain.
@@
The remarkably good performance of formula (12) in describing the numerically computed $\lambda_{\text{p}}(L)$ is demostrated in Fig.4. The one can observe that it gives an excellent quantitative description also in the regime of large differences between $\lambda_{\text{p}}(L)$ and $\lambda_{\text{p,max}}$. In this large-deviations regime, the formula (12) has to be further supplemented by the condition that, if it predicts $\lambda_{\text{p}}<0$, then $\lambda_{\text{p}}=0$ should be substituted instead, implying that the periodic motion is Lyapunov stable. With such a modification, formula (12) turns into an excellent predictor of the lengths $L$ corresponding to the stable periodic motion. We, finally, note, that the quality of the approximation (12) can be improved further by adding there terms with higher-order powers of $(q_{0}-q_{\text{p}})$.
@@
The function $\lambda_{\text{p}}(J)$ shown in Fig. 3 exhibits tendency to diverge on the approach of $J$ to the numerically computed separatrix value $J^{*}=1.1504059085$. Yet, due to the finiteness of the Hamiltonian parameters $J$ and $h$ and the smoothness of the dynamics, the rate of the divergence between two close phase space trajectories is supposed to be limited from above by a number of the order of $J$ or $h$, hence the value of $\lambda_{\text{p}}(J)$ cannot diverge.
In this part, we zoom in on the behavior $\lambda_{\text{p}}(J)$ in the close vicinity of the separatrix shown in Fig. 2. As schematically illustrated in Fig. 6, the separatrix consists of an unstable fixed point analogous to an upward pointing pendulum and two trajectories starting infinitesimally close to that fixed point at time $t=-\infty$ and approaching it infinitesimally close at $t=+\infty$. The above fixed point is simultaneously a many-spin trajectory belonging to the $q=0$ subspace of the full phase space. This trajectory is unique in the sense that, in comparison with the previously considered periodic trajectories, it possesses an additional instability (towards the branches of the separatrix) developing entirely within the $q=0$ subspace, as opposed to involving the subspaces with $q\neq 0$. In the case considered below, that additional instability actually has a Lyapunov exponent which is larger than the largest one for the periodic trajectories. The results of our numerical calculation of $\lambda_{\text{p}}(J)$ for a chain of 10 spins in the close vicinity of the separatrix value $J=J^{*}$ are presented Fig. 7. This calculation becomes increasingly demanding computationally as $J$ approaches $J^{*}$, because (i) the topological character of the periodic trajectories changes from a connected libration to two disconnected rotations, and (ii) this change is accompanied by the slowing down of the dynamics. Our findings indicate that that $\lambda_{\text{p}}(J)$ for $J\to J^{*}$ has the form of a symmetric cusp approaching the finite value $\lambda_{A}=1.6$ with infinite derivative $\lambda^{\prime}_{\text{p}}(J)$:
@@
Yet, as indicated in Fig. 7, for the exact separatrix value $J=J^{*}$, we obtain another Lyapunov exponent $\lambda_{\text{p}}(J^{*})\equiv\lambda_{S}=1.99>\lambda_{A}$. The latter result is computed by choosing the unstable fixed point on the separatrix as the reference trajectory $\mathcal{S}(t)$ and then extracting $\lambda_{S}$ as the exponential growth rate of the small deviations $|\delta\mathcal{S}(t)|$ starting from $|\delta\mathcal{S}(0)|\approx 10^{-6}\sqrt{L}$ and letting it grow up to $10^{-1}\sqrt{L}$ without making any resets. In a related finding, we also observed switching of the wave number $q_{\text{p}}(J)$ corresponding to $\lambda_{\text{p}}(J)$, namely: $q_{\text{p}}(J^{*})=0$, while $q_{\text{p}}(J)=\pi/5$ for $J\neq J^{*}$ in the vicinity of the separatrix.
@@
where $\Delta J\equiv J-J^{*}$, $\lambda_{A}$ is the asymptotic value of $\lambda_{\text{p}}$ defined by Eq.(13) and $C$ is a constant. The derivation of Eq.(15) is relegated to Appendix D. Here we only mention that Fig. 7 demonstrates that the functional form (15) is quantitatively consistent with our numerical results.
@@
In this part, we propose a method of experimentally observing the Lyapunov instability around a periodic trajectory in a system of many classical spins. The challenge here is that a direct method based on a very accurate control and monitoring of system’s phase trajectory in many-particle phase space is, likely, impractical. As we show below, it is, in fact, not necessary to monitor the entire phase space trajectory – rather it is sufficient to monitor just one observable. A similar approach was adopted previously in Refs. [59, 60, 123, 45], where the Lyapunov exponents for the ergodic trajectories were obtained through monitoring ensemble-averaged time evolution of a single variable in the course of a slightly imperfect time reversal of the dynamics. [In that case, the approach to extracting the Lyapunov exponents involves “out-of-time-order correlators”(OTOCs)[63].]

### v3

@@ IV.1 Equations of motion
Where necessary, the set of equations (6, 7) is to be solved numerically using the 4th-order Runge-Kutta algorithm with the discretization time steps $\delta t$ being small enough to make the results insensitive to making $\delta t$ smaller; typically $\delta t=0.001$.
@@
where the projections are the solutions of Eqs.(6, 7).
@@ IV.2 Periodic trajectories
Due to the translational invariance of both the Hamiltonian (III) and the initial conditions (3), the classical spins initially polarised in the same direction exhibit the same time evolutions and hence remain parallel to each other at all subsequent moments of time with the direction of their common polarisation evolving as a function of time. As a result, the calculation of a many-spin trajectory $\mathcal{S}(t)$ is reduced to computing the same one-spin trajectory $\mathbf{S}_{i}(t)=\mathbf{S}_{\text{p}}(t)$ for each spin on the lattice. Since the one-spin trajectory $\mathbf{S}_{\text{p}}(t)$ is limited to a two-dimensional unit sphere $|\mathbf{S_{\text{p}}}|=1$, it normally closes onto itself and thus necessarily becomes periodic. (This property is explicitly demonstrated for our system in Appendix A. The broader context is that of the Poincaré–Bendixson theorem [127, 128].) The many-spin phase space trajectory $\mathcal{S}_{\text{p}}(t)$ corresponding to $\mathbf{S}_{\text{p}}(t)$ also becomes periodic.
@@ IV.3 Librations and rotations
Dependent on the parameters of the Hamiltonian $\mathcal{H}_{\text{p}}$ and on the initial orientation $\mathbf{S}_{\text{p}}(0)$, there are two possible kinds of generic periodic trajectories $\mathbf{S}_{\text{p}}(t)$, which, by analogy with the description of a pendulum[43], we call “librations” and “rotations”. Librations are realised by a single trajectory on the spherical surface $|\mathbf{S_{\text{p}}}|=1$ connecting all points having the same energy. The rotations are realised by two disconnected trajectories – both corresponding to the same energy. Librations and rotations transition to each other as a function of Hamiltonian parameters and/or initial energy through a separatrix as illustrated in Fig.2. (See Appendix A for an explicit derivation.)
@@ IV.4 Lyapunov instabilities: general considerations
The finiteness of $\lambda_{\text{max}}$ for the lattices of classical spins with finite number of interacting neighbors is proven in Appendix B following the general argument of Ref. [126]. Practical numerical computation of $\lambda_{\text{max}}$ requires one to perform a large number of resets (contractions) of $|\delta\mathcal{S}(t)|$ as described in Refs.[129, 130, 131] and in Appendix C.
Largest Lyapunov exponents and the entire Lyapunov spectra of classical spin systems have previously been investigated [44, 126] but only in the vicinity of ergodic trajectories (also called “stochastic”), which nonperiodically cover the energy shell of the system. If a trajectory is ergodic, it passes in the vicinity of every point on the ergodic part of the energy shell, thereby “collecting” all possible local growth rates of $|\delta\mathcal{S}(t)|$ and averaging over them. As a result, the largest “ergodic” Lyapunov exponent, to be denoted as $\lambda_{\text{e}}$, and the entire Lyapunov spectrum do not depend on the choice of the initial conditions $|\mathcal{S}(0)|$ [129]. The convergence of the procedure for calculating $\lambda_{\text{e}}$ is, in fact, a good measure of the ergodisation time of the underlying dynamics[45]. Plots of $\lambda_{\text{e}}$ computed for the ergodic trajectories of our spin chain with different $J$ and $L$ can be found in Appendix D.
@@
Generically, different wave vectors $\mathbf{q}_{k}$ are supposed to correspond to different values of $\lambda_{k}$ with the exception of symmetry-related or accidental degeneracies. A symmetry-related degeneracy occurs, in particular, when a lattice, such as the one we consider, has inversion symmetry, and, as a result, pairs of wave vectors ±$\pm\mathbf{q}_{k}$ correspond to the same $\lambda_{k}$. In this case, a generic Lyapunov exponent is at least double-degenerate, the possible non-degenerate cases correspond to $\mathbf{q}_{k}=0$ and $\mathbf{q}_{k}=\pi/L$.
@@ IV.5.1 Overview
Figure 4: Dependence of periodic Lyapunov exponent $\lambda_{\text{p}}$ on chain length $L$ for three different values of $J$: (a)  $J=0.79$, (b) $J=1.76$ and (c) $J=2.23$. Panel (a) corresponds to librations, while panels (b) and (c) correspond to rotations. Points represent numerically computed $\lambda_{\text{p}}$. Red lines are the fits based on Eq.(13) with parameters indicated in the plot legends. Points marked as “Stable” correspond to the $\lambda_{\text{p}}=0$ verified as explained in Appendix C. In a few cases marked as entering Arnold diffusion regime, Lyapunov instability around a periodic trajectory is followed by anomalously long-living nearly quasiperiodic behavior — see Section V for further discussion.
@@
Overall, the strong and nontrivial $J$- and $L$-dependencies of the largest Lyapunov exponent $\lambda_{\text{p}}$ for periodic trajectories (Figs. 3 and 4) should be compared with the rather featureless $J$- and $L$-dependencies of the largest Lyapunov exponent $\lambda_{\text{e}}$ for nonperiodic ergodic trajectories on the same energy shell (Fig. 24 in Appendix D). Let us also mention that $\lambda_{\text{p}}$ can be either larger or smaller than $\lambda_{\text{e}}$. For example, $\lambda_{\text{p}}\approx 0.35$ for a chain with $J=1.76$ in the limit of large $L$ [see Fig. 4(b)], which is significantly smaller than the ergodic exponent $\lambda_{\text{e}}\approx 1.13$ for the same $J$ (see Fig. 24). On the other hand, the periodic exponent near the separatrix $J=J^{*}$ reaches the value $\lambda_{\text{p}}\simeq 1.43$ (see Fig. 7), which is noticeably larger than the ergodic exponent $\lambda_{\text{e}}\simeq 0.67$ in Fig. 24 for the same $J$.
@@
Figure 5: Spatial Fourier spectra $f_{\lambda_{\text{p}}}(q)$ of the Lyapunov vectors defined by Eq. (12). The chain lengths $L$ and the coupling constants $J$ are indicated in plot legends. The allowed values of wavenumbers are $q=\frac{2\pi}{L}k$, where $k$ is an integer in the interval $[-L/2,L/2]$, $t_{0}=750$.
@@
$f_{\lambda_{\text{p}}}(q,t)\equiv\frac{1}{L}\left|\sum_{m=1}^{m=L}\delta\mathbf{S}_{m}(t)\ e^{-iqm}\right|^{2}$ (12)
of the Lyapunov vectors $\delta{\cal S}(t)$ corresponding to $\lambda_{\text{p}}$ with time $t$ being large enough for the computed $\lambda_{\text{p}}$ to reach the stationary value. Here $\delta\mathbf{S}_{m}\equiv(\delta S_{m}^{x},\delta S_{m}^{y},\delta S_{m}^{z})$ are the components of the high-dimensional vector $\delta{\cal S}$ entering Eq.(11); the Fourier amplitude inside the absolute-value bars |||...|| … | is a three-dimensional vector with $x$, $y$ and $z$ complex-valued projections, $|...|^{2}$ is the sum of the squares of the moduli of those projections. The Fourier spectra in Fig. 5 are computed for two chain lengths $L=18$ and $L=20$ and for two interaction constants $J$ representing the libration and the rotation regimes for each length. In all four cases, one can see two dominant Fourier peaks corresponding to ±$\pm q_{\text{p}}$.
The underlying translational symmetry breaking helps us to explain the distinct behavior of $\lambda_{\text{p}}(J)$ plotted in Fig.3: In both panels of Fig.3, there are intervals of $J$ where Lyapunov exponents $\lambda_{\text{p}}$ for two chains of different lengths coincide with each other, and also there are intervals where the two Lyapunov exponents suddenly become different, because one of of the two functions $\lambda_{\text{p}}(J)$ exhibits a kink. The pair of plots in Fig. 3(a) and another pair in Fig. 3(b) – both correspond to chain lengths such that one chain is twice the length of the other. Therefore, the wave numbers $q=\frac{2\pi}{L}n$, where $n$ is an integer and $L$ the length of the shorter chain, are the allowed wave numbers for the both chains, but then the longer chain has additional allowed wave numbers that are not available for the shorter one. Thus, if the largest Lyapunov exponent for the longer chain corresponds to $q$ available for the shorter chain, then the two Lyapunov exponents $\lambda_{\text{p}}$ are equal to each other. If, however, the above $q$ for the longer chain is not available for the shorter one, then the values of $\lambda_{\text{p}}$ become different, with the larger $\lambda_{\text{p}}$ necessarily corresponding to the longer chain. The fact that the above kind of switching behavior is exhibited in Fig. 3 only in the rotation regime is, probably, a coincidence; we did not identify any fundamental reason for the librations not to exhibit such a switching.
@@
The remarkably good performance of formula (13) in describing the numerically computed $\lambda_{\text{p}}(L)$ is demostrated in Fig.4. One can observe that it gives an excellent quantitative description also in the regime of large differences between $\lambda_{\text{p}}(L)$ and $\lambda_{\text{p,max}}$. In this large-deviations regime, the formula (13) has to be further supplemented by the condition that, if it predicts $\lambda_{\text{p}}<0$, then $\lambda_{\text{p}}=0$ should be substituted instead, implying that the periodic motion is Lyapunov stable. With such a modification, formula (13) turns into an excellent predictor of the lengths $L$ corresponding to the stable periodic motion. We, finally, note, that the quality of the approximation (13) can be improved further by adding there terms with higher-order powers of $(q_{0}-q_{\text{p}})$.
@@
The function $\lambda_{\text{p}}(J)$ shown in Fig. 3 exhibits tendency to diverge on the approach of $J$ to the numerically computed separatrix value $J^{*}=1.1504059085$. Yet, due to the finiteness of the Hamiltonian parameters $J$ and $h$ and the smoothness of the dynamics, the rate of the divergence between two close phase space trajectories is limited from above by a number of the order of $J$ or $h$, hence the value of $\lambda_{\text{p}}(J)$ cannot diverge (see Appendix B for the derivation).
In this part, we zoom in on the behavior $\lambda_{\text{p}}(J)$ in the close vicinity of the separatrix shown in Fig. 2. As schematically illustrated in Fig. 6, the separatrix consists of an unstable fixed point analogous to an upward pointing pendulum and two trajectories starting infinitesimally close to that fixed point at time $t=-\infty$ and approaching it infinitesimally close at $t=+\infty$. The above fixed point is simultaneously a many-spin trajectory belonging to the $q=0$ subspace of the full phase space. This trajectory is unique in the sense that, in comparison with the previously considered periodic trajectories, it possesses an additional instability (towards the branches of the separatrix) developing entirely within the $q=0$ subspace, as opposed to involving the subspaces with $q\neq 0$. In the case considered below, that additional instability actually has a Lyapunov exponent which is larger than the largest one for the periodic trajectories. The results of our numerical calculation of $\lambda_{\text{p}}(J)$ for a chain of 10 spins in the close vicinity of the separatrix value $J=J^{*}$ are presented Fig. 7. This calculation becomes increasingly demanding computationally as $J$ approaches $J^{*}$, because (i) the topological character of the periodic trajectories changes from a connected libration to two disconnected rotations, and (ii) this change is accompanied by the slowing down of the dynamics. Our findings indicate that $\lambda_{\text{p}}(J)$ for $J\to J^{*}$ has the form of a symmetric cusp approaching the finite value $\lambda_{A}=1.6$ with infinite derivative $\lambda^{\prime}_{\text{p}}(J)$:
@@
Yet, as indicated in Fig. 7, for the exact separatrix value $J=J^{*}$, we obtain another Lyapunov exponent $\lambda_{\text{p}}(J^{*})\equiv\lambda_{S}=1.99>\lambda_{A}$. The latter result is computed by choosing the unstable fixed point on the separatrix as the reference trajectory $\mathcal{S}(t)$ and then extracting $\lambda_{S}$ as the exponential growth rate of the small deviations $|\delta\mathcal{S}(t)|$ starting from $|\delta\mathcal{S}(0)|\approx 10^{-6}\sqrt{L}$ and letting it grow up to $10^{-1}\sqrt{L}$ without making any resets (see Appendix E for further details). In a related finding, we also observed a switching of the wave number $q_{\text{p}}(J)$ corresponding to $\lambda_{\text{p}}(J)$, namely: $q_{\text{p}}(J^{*})=0$, while $q_{\text{p}}(J)=\pi/5$ for $J\neq J^{*}$ in the vicinity of the separatrix.
@@
where $\Delta J\equiv J-J^{*}$, $\lambda_{A}$ is the asymptotic value of $\lambda_{\text{p}}$ defined by Eq.(14) and $C$ is a constant. The derivation of Eq.(16) is relegated to Appendix F. Here we only mention that Fig. 7 demonstrates that the functional form (16) is quantitatively consistent with our numerical results.
@@
In this part, we propose a method of experimentally observing the Lyapunov instability around a periodic trajectory in a system of many classical spins. A possible realization may be based on mechanical spin-like simulators, such as those implemented in Ref. [132]. The challenge here is that a direct method based on a very accurate control and monitoring of system’s phase trajectory in many-particle phase space is, likely, impractical. As we show below, it is, in fact, not necessary to monitor the entire phase space trajectory – rather it is sufficient to monitor just one observable. A similar approach was adopted previously in Refs. [59, 60, 133, 45], where the Lyapunov exponents for the ergodic trajectories were obtained through monitoring ensemble-averaged time evolution of a single variable in the course of a slightly imperfect time reversal of the dynamics. [In that case, the approach to extracting the Lyapunov exponents involves “out-of-time-order correlators”(OTOCs)[63].]